Phantomdex:

The corn crop yield based on a new fertilizer is modeled by the equation ŷ = 0.01x + 1.24, where x represents the amount of fertilizer in pounds per acre and ŷ represents the predicted yield of crops in tons per acre. How much fertilizer should be used to yield 7 tons of corn per acre? 1.31 pounds per acre 5.76 pounds per acre 131 pounds per acre 576 pounds per acre

Midnight97:

Alright, let's see here

Phantomdex:

I have like less than 10 minutes left

Midnight97:

@phantomdex wrote:
I have like less than 10 minutes left
Ik I am trying to be fast

Midnight97:

y=0.01x+1.24 7=0.01x+1.24 7-1.24=0.01x 5.76=0.01x x=5.76/0.01

Midnight97:

576 pounds per acre is your answer